Assume that adults have scores that are normally distributed with a mean of μ-105 and a standard dev ation ơt5 Find the probatify that a randomly selected adult has an IQ less than 132 Click to view page 1 of the table Click to view page 2 of the table The probability that a randomly selected adult has an IQ less than 132 is (Type an integer or decimal rounded to four decimal places as needed)
Answer #1

Solution:

Given in the question

Mean = 105

Standard deviation = 15

P(Xbar< 132)

z =(132-105)/15 = 27/15 = 1.8

So from z table we found p- value

P(Xbar<132) = 0.9641

So there is 96.41% probability that IQ is less than 132.